OpenStudy (anonymous):

37=-3+5(x+6) help please

OpenStudy (anonymous):

37= -3 + 5x + 30 10 = 5x x = 2

jimthompson5910 (jim_thompson5910):

37=-3+5(x+6) 37=-3+5x+30 37=5x+27 37-27=5x 10=5x 10/5=x 2=x x=2

OpenStudy (anonymous):

ok so on the step 37=5x+27 why didnt you do anthing to the three??

jimthompson5910 (jim_thompson5910):

-3+30 = 27 that's where the -3 went